OpenStudy (anonymous):

What is the quadratic function that is created with roots -3 and 1 and a vertex at (-1, -8)?

OpenStudy (anonymous):

in root form (if there is such a thing), you'd write it like this... \[f(x)=a(x-r_1)(x-r_2)\]where \(r_1 \text{ & } r_2\) are the roots and \(\text{a}\) is a constant. to find the value of the constant, simply plug in the values for the roots and the given point, in this case the vertex, and solve for a.

OpenStudy (anonymous):

So f(x) = a(-1+3)(-1+8) f(x) = a(2)(7) f(x) = 14a Where do I go from there?

OpenStudy (anonymous):

\[f(-1) = y = -8\] so it should be like this...\[f(-1)=a(-1+3)(-1-1) \Rightarrow -8 = a(2)(-2) \Rightarrow -8 =-4a \Rightarrow a=2\]